OpenStudy (anonymous):

I WILL FAN AND MEDAL PLEASE HELP!!! A doctor is doing a study of the height of children who eat their vegetables. He recorded the height of the children each year. The first year, the child was 36 inches. The second year, the child was 40 inches. The third year, the child was 44 inches. Write a function rule to represent the height of the child as an arithmetic sequence. How tall will the child be after 6 years?

OpenStudy (anonymous):

a1=36in a2=40in a3=44in do u see the relation between each one and the next one?

OpenStudy (anonymous):

Are there options? If not, the kid grows 4 inches a year, so 44+4 = 48 (fourth year) 48+4=52 (fifth year) 52+4=56 (sixth year)

OpenStudy (anonymous):

@kaylie.sullivan is right

OpenStudy (anonymous):

Basically 4 inches each year

OpenStudy (anonymous):

yes I do now and its not multiple choice but thank you it just wouldn't click for me at first I guess

OpenStudy (anonymous):

np girl

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

:)